# Cross-modal Cognitive Consensus guided Audio-Visual Segmentation

### v4

@@ Abstract
Audio-Visual Segmentation (AVS) aims to extract the sounding object from a video frame, which is represented by a pixel-wise segmentation mask for application scenarios such as multi-modal video editing, augmented reality, and intelligent robot systems. The pioneering work conducts this task through dense feature-level audio-visual interaction, which ignores the dimension gap between different modalities. More specifically, the audio clip could only provide a Global semantic label in each sequence, but the video frame covers multiple semantic objects across different Local regions, which leads to mislocalization of the representationally similar but semantically different object. In this paper, we propose a Cross-modal Cognitive Consensus guided Network (C3N) to align the audio-visual semantics from the global dimension and progressively inject them into the local regions via an attention mechanism. Firstly, a Cross-modal Cognitive Consensus Inference Module (C3IM) is developed to extract a unified-modal label by integrating audio/visual classification confidence and similarities of modality-agnostic label embeddings. Then, we feed the unified-modal label back to the visual backbone as the explicit semantic-level guidance via a Cognitive Consensus guided Attention Module (CCAM), which highlights the local features corresponding to the interested object. Extensive experiments on the Single Sound Source Segmentation (S4) setting and Multiple Sound Source Segmentation (MS3) setting of the AVSBench dataset demonstrate the effectiveness of the proposed method, which achieves state-of-the-art performance. Code will be available at https://github.com/ZhaofengSHI/AVS-C3N once accepted.
@@ II-A Visual Object Segmentation
II-B Sound Source Localization and Segmentation

### v5

@@ Abstract
Audio-Visual Segmentation (AVS) aims to extract the sounding object from a video frame, which is represented by a pixel-wise segmentation mask for application scenarios such as multi-modal video editing, augmented reality, and intelligent robot systems. The pioneering work conducts this task through dense feature-level audio-visual interaction, which ignores the dimension gap between different modalities. More specifically, the audio clip could only provide a Global semantic label in each sequence, but the video frame covers multiple semantic objects across different Local regions, which leads to mislocalization of the representationally similar but semantically different object. In this paper, we propose a Cross-modal Cognitive Consensus guided Network (C3N) to align the audio-visual semantics from the global dimension and progressively inject them into the local regions via an attention mechanism. Firstly, a Cross-modal Cognitive Consensus Inference Module (C3IM) is developed to extract a unified-modal label by integrating audio/visual classification confidence and similarities of modality-agnostic label embeddings. Then, we feed the unified-modal label back to the visual backbone as the explicit semantic-level guidance via a Cognitive Consensus guided Attention Module (CCAM), which highlights the local features corresponding to the interested object. Extensive experiments on the Single Sound Source Segmentation (S4) setting and Multiple Sound Source Segmentation (MS3) setting of the AVSBench dataset demonstrate the effectiveness of the proposed method, which achieves state-of-the-art performance. Code is available at https://github.com/ZhaofengSHI/AVS-C3N.
@@ II-A Visual Object Segmentation
II-B Referring Video Object Segmentation
Referring Video Object Segmentation (RVOS) aims to extract pixel-level segmentation masks of the interesting objects in the video according to the given language descriptions. Gavrilyuk et al. [57] make the first attempt towards the RVOS task. URVOS [58] constructs a large-scale dataset and proposes a unified model, which utilizes the mask propagation operation. VTCapsule [59] proposes a capsule-based method for learning effective representations. Some methods [60, 61] integrate various attention mechanisms into RVOS frameworks to enhance the visual and text features. Ding et al. [62, 63], Feng et al. [64], and Yang et al. [65] introduce hierarchical visual-language feature fusion methods. MTTR [66] and ReferFormer [67] model the task as a sequence prediction problem and propose Transformer-based networks to simplify the RVOS pipeline. Recently, Lan et al. [68] propose the awesome BIFIT framework, which performs inter-frame interaction and enhances the bilateral correlations between the linguistic and visual features. Luo et al. [69] aggregate the semantic-level visual and textual information for temporal modeling and cross-modal alignment. Miao et al. [70] use the spectrum-domain information for performing global interaction and extracting effective multimodal representations.
Although the RVOS task introduces an additional text modality, it requires user interaction to provide linguistic guidance, whereas the AVS task leverages the inherent audio information from the input videos.
II-C Sound Source Localization and Segmentation
@@ IV-D8 Evaluation under varying visual qualities
IV-D9 Comparison between C3IM and other cross-modal interaction modules
| Modules | mIoU | F-score | ||
| S4 | MS3 | S4 | MS3 | |
| C3IM | 82.94 | 61.67 | 0.906 | 0.713 |
| MHA [120, 121] | 82.16 | 61.20 | 0.901 | 0.706 |
| BFM [122] | 81.46 | 59.50 | 0.899 | 0.701 |
| TPAVI [14] | 80.81 | 58.69 | 0.903 | 0.684 |
To further demonstrate the superiority of the proposed C3IM, we conduct detailed comparative experiments that replace the C3IM in our C3N (PVTv2+PANNs) with other cross-model interaction modules. In detail, we choose three recent cross-modal multi-head attention (MHA) module used in [120, 121], Bilateral-Fusion Module (BFM) used in [122], and the temporal pixel-wise audio-visual interaction (TPAVI) module used in [14] as alternatives to the C3IM. Experimental results in Table VIII show our C3IM outperforms other powerful cross-model interaction modules under the S4 and MS3 settings. Specifically, our C3IM achieves higher mIoU than the MHA, BFM, and TPAVI by 0.78, 1.48, and 2.13, respectively. Under the MS3 setting, the C3IM surpasses the MHA, BFM, and TPAVI by 0.47, 2.17, and 2.98 in mIoU, and 0.007, 0.012, and 0.029 in F-score, respectively.
@@ V Conclusion
For the AVS task, the sounding objects in each frame of the video are frequently temporally correlated. Thus, it is also necessary to design inter-frame interaction methods to enhance the frame-to-frame correlations of the segmented objects. The design of the inter-frame interaction method needs to focus on two aspects: On the one hand, when the appearance of the sounding object suddenly changes such as it is highly occluded, the inter-frame compensational information should be utilized to accurately identify and segment the sounding object. On the other hand, when the audio abruptly changes, the model needs to capture the inter-frame discrepancy information to exclude the irrelevant objects in the video. In the future, we will further explore the modeling of the frame-to-frame correlations of the sounding objects.
Acknowledgments
This work was supported in part by the National Science and Technology Major Project (2021ZD0112001), the National Natural Science Foundation of China (No. U23A20286), the Independent Research Project of Civil Aviation Flight Technology and Flight Safety Key Laboratory (FZ2022ZZ06), and the Natural Science Foundation of Sichuan Province (2023NSFSC1972).